### v1

# Feshbach-Villars Formalism for a Spin-1/2 Particle in Curved Spacetime

Feshbach-Villars Formalism for a Spin-1/2 Particle in Curved Spacetime
Abdelmalek Boumali abdelmalek.boumali@univ-tebessa.dz or boumali.abdelmalek@gmail.com Laboratory of theoretical and applied Physics,
Echahid Cheikh Larbi Tebessi University, Algeria
(April 1, 2025)
This study explores the Feshbach-Villars (FV) formalism for spin-1/2 particles in curved spacetime. We derive the Hamiltonian form of the Dirac equation in this context and extend the FV transformation accordingly. The generalized Klein-Gordon equation, obtained by squaring the Dirac operator, is reformulated using the FV approach. We present the resulting Hamiltonian in both matrix and Pauli matrix forms, considering gravitational and electromagnetic interactions. The formalism is examined in both (1+2) and (1+3) dimensional spacetimes, with special attention given to the spin-field interaction term. This study provides a framework for studying relativistic quantum mechanics in curved spacetime, offering insights into the interplay between quantum effects, gravity, and electromagnetism.
I Introduction
Relativistic quantum mechanics represents a significant academic achievement that focuses on integrating the concepts of relativity and quantum mechanics. This integration expanded the framework of quantum theory and introduced the concept of negative energy levels, which emerged directly from relativistic formulations. Negative energy states are illustrated by the concept of the Dirac sea, which is an infinite expanse of particles with negative energy. In this model, the absence of a particle in the Dirac sea indicates the presence of a physical particle, specifically, a positron. This interpretation arises from the requirement that the energy of free particles must be positive, as dictated by the physical laws. By treating particles and antiparticles differently, this approach creates intrinsic asymmetry. While this asymmetry is essential for upholding the probabilistic principles of quantum physics, as outlined in the continuity equation, it also highlights the elegance of Dirac’s sea model. This concept is particularly applicable to fermions, which have a half-integer spin and obey the Pauli exclusion principle. However, this does not apply to bosons, which possess an integer spin. Quantum field theory (QFT) provides a more comprehensive framework by treating antiparticles as fundamental components of field interactions rather than relying on the Dirac sea [1, 2, 3].
In 1958, Feshbach and Villars [1] published their findings on linearizing the Klein-Gordon equation. Their work led to the development of the Feshbach-Villars (FV-0) equations, which allowed the Klein-Gordon equation to be expressed in the Schrödinger form, also known as the Hamiltonian form. This formalization addresses various contradictions in the Klein-Gordon equation, including the existence of negative energies. This facilitated a clearer interpretation of probability theory, particularly when considering particles with corresponding antiparticles.
Feshbach and Villars’ formalism resolved the key issues that had previously caused the abandonment of the Klein-Gordon theory, enabling a reevaluation of its principles. Significant research has been conducted using the Feshbach-Villars equations. Furthermore, efforts have been made to obtain numerical solutions and approximations for these equations, providing physicists with the means to tackle ongoing challenges and develop insights related to relativistic quantum mechanics (RQM) [4, 5, 6, 7, 8, 9, 10].
To expand the theory within the framework of general relativity, Silenko [11, 12] was the first to introduce a new form of conformal invariance for massless scalar particles in Riemannian spacetimes. He studied scalar particles in arbitrary spacetimes under general relativity (GR). To derive a Hamiltonian form of the covariant Klein-Gordon (KG) equation for both massive and massless particles, he utilized the generalized Feshbach-Villars transformation, as proposed in a previous study. Silenko then derived the general quantum-mechanical equations of motion and established their classical limits. Recently, this method has been applied in various contexts, and the results have been promising [13, 14, 15, 16].
The Feshbach-Villars equations have simplified the reevaluation of relativistic quantum mechanics for spin-0 particles and opened new avenues for investigating similar theories related to spin-1/2 particles. In addition, this formalism offers an alternative approach for spin-zero particles characterized by the Klein-Gordon equation (KG-0). This method incorporates antiparticles through charge density instead of probability density, thereby expanding the representation space to a pseudo-Hilbert space. Notably, this framework treats particles and antiparticles symmetrically, eliminating the need for a Dirac sea and enhancing the mathematical representation. This has significant implications for our understanding of the spin-zero particles and their interactions.
Building on this idea, FV-like equations have been proposed for spin-½ particles (KG–½). These equations have the potential to achieve complete particle-antiparticle symmetry and eliminate the need for a Dirac sea, representing a promising development. By altering the manner in which wave functions depict physical states, this extension alters certain scientific interpretations, particularly those involving transition probabilities. The differences become significant when examining phenomena where the influence of the Dirac sea is evident. The formalism presented by Feshbach and Villars paved the way for a new relativistic quantum mechanics equation for spin-1/2 particles. The worked developed by Robson and Staudte [2, 3] followed the idea of the Feshbach-Villars linearization procedure and applied it to the second-order Dirac equation. The result was an eight-component relativistic equation in Hamiltonian form, which preserves some of the inner product properties and the probabilistic interpretation obtained from the FV-0 equation. This eight-component equation or Feshbach-Villars equation for spin- 1 / 2 particles (FV-1/2) is relativistic and different from the Dirac equation.
This study aims to extend the Feshbach-Villars (FV) formalism for spin-1/2 particles to curved spacetime. Specifically, we analyze fermion particles in arbitrary spacetimes within the framework of general relativity (GR). To derive the Hamiltonian form of the covariant Dirac equation in curved spacetime, we employ the generalized Feshbach-Villars transformation introduced in Ref. [12]. It is worth noting that, to the best of our knowledge, there is a notable gap in the literature regarding this type of study.
II Hamiltonian form of Dirac equation in curved spacetime
The Hamiltonian form of the Dirac equation in curved spacetime involves recasting the covariant form of the equation into a time-evolution operator that acts on the wavefunction.
The Dirac equation in curved spacetime is:
$\left[i\gamma^{\mu}\mathcal{D}_{\mu}-m\right]\Psi=0,$ (1)
where:
  • $\gamma^{\mu}=e_{a}^{\mu}\gamma^{a}$are the curved spacetime gamma matrices, with $e_{a}^{\mu}$ being the tetrads and $\gamma^{a}$ the flat spacetime gamma matrices.
  • $\mathcal{D}_{\mu}=\partial_{\mu}+\frac{1}{4}\omega_{\mu}^{ab}\gamma_{ab}$ is the covariant derivative, including the spin connection $\omega_{\mu}^{ab}$.
  • $m$ is the mass of the spinor field.
Expanding, and rearranging to isolate the time derivative we have
$i\gamma^{0}\mathcal{D}_{0}\Psi=\left[-i\gamma^{i}\mathcal{D}_{i}+m\right]\Psi.$ (2)
wher
$\mathcal{D}_{0}=\partial_{0}+\frac{1}{4}\omega_{0}^{ab}\gamma_{ab},\,\mathcal{D}_{i}=\partial_{i}+\frac{1}{4}\omega_{i}^{ab}\gamma_{ab}$ (3)
. (4)
Multiply (2) through by$(\gamma^{0})^{-1}$ (note: in most conventions, $(\gamma^{0})^{-1}=\gamma^{0}$), give
$i\partial_{0}\Psi=\gamma^{0}\left[-i\gamma^{i}\left(\partial_{i}+\frac{1}{4}\omega_{i}^{ab}\gamma_{ab}\right)+m-\frac{1}{4}\omega_{0}^{ab}\gamma_{ab}\right]\Psi.$ (5)
This gives the Hamiltonian form:
$i\partial_{0}\Psi=\mathcal{H}\Psi,$ (6)
where
$\mathcal{H}=\gamma^{0}\left[-i\gamma^{i}\mathcal{D}_{i}+m-\frac{1}{4}\omega_{0}^{ab}\gamma_{ab}\right].$ (7)
Here
  • $-i\gamma^{0}\gamma^{i}\mathcal{D}_{i}$: Represents the kinetic energy and coupling of the spinor to the spatial curvature via the spin connection $\omega_{i}^{ab}$.
  • $m\gamma^{0}$: Represents the mass term in the Hamiltonian.
  • $-\frac{1}{4}\gamma^{0}\omega_{0}^{ab}\gamma_{ab}$: Represents the coupling to the spin connection in the time direction.
Final Form of the Hamiltonian
$\mathcal{H}=\gamma^{0}\left[-i\gamma^{i}\left(\partial_{i}+\frac{1}{4}\omega_{i}^{ab}\gamma_{ab}\right)+m-\frac{1}{4}\omega_{0}^{ab}\gamma_{ab}\right].$ (8)
This Hamiltonian governs the time evolution of the spinor field $\Psi$ in curved spacetime, accounting for both gravitational and gauge interactions.
The probability density for the Dirac equation in curved spacetime is a generalization of the flat spacetime case, reflecting the local geometry and the covariant nature of the wavefunction. In curved spacetime, the probability density is closely tied to the conserved current derived from the Dirac equation. These quantities is defined as the time component of the conserved current $J^{\mu}$, where:
$J^{\mu}=\bar{\psi}\gamma^{\mu}\psi,$ (9)
and $\bar{\psi}=\psi^{\dagger}\gamma^{0}$ is the Dirac adjoint. The conservation of the current $J^{\mu}$ is guaranteed by the Dirac equation:
$\nabla_{\mu}J^{\mu}=0,$ (10)
where $\nabla_{\mu}$ is the covariant derivative compatible with the curved spacetime metric. Thus, the probability density $\rho$ is the 00-component of the current:
$\rho=J^{0}=\bar{\psi}\gamma^{0}\psi.$ (11)
In curved spacetime, the interpretation of $\rho$ is influenced by the metric determinant $\sqrt{-g}$, which ensures proper normalization. To properly normalize the wavefunction in curved spacetime, we integrate $\rho$ over a spacelike hypersurface $\Sigma$:
$\int_{\Sigma}\rho\,\sqrt{-g}\,\text{d}^{3}x=1,$ (12)
where $\sqrt{-g}$ is the determinant of the spacetime metric $g_{\mu\nu}$, ensuring the proper volume element in curved spacetime. Despite the curvature, $J^{\mu}$ is conserved globally due to $\nabla_{\mu}J^{\mu}=0$, ensuring that the total probability remains constant.
III Feshbash-Villars of a particle of spin-1/2 in curved spacetime: The formalism
The modified Klein-Gordon equation obtained by squaring the operator in the Dirac equation, first found by Erwin Schrödinger is given by
$\left(\frac{1}{\sqrt{-g}}\mathcal{D}_{\mu}\left(\sqrt{-g}g^{\mu\nu}\mathcal{D}_{\nu}\right)-\frac{1}{4}R+\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}+m^{2}\right)\Psi=0$ (13)
where $R$ is the Ricci scalar, and $F_{\mu\nu}$ is the field strength of $A_{\mu}$ and
$\mathcal{D}_{\mu}=\partial_{\mu}+\frac{1}{4}\omega_{\mu}^{ab}\gamma_{ab}-iqA_{\mu}$ (14)
Following the procedure used by Silenko (12, 2025) the general form of the covariant KG equation reads(13) (see Annex A)
$\left(\partial_{0}^{2}+\frac{1}{g^{00}\sqrt{-g}}\left\{\partial_{i},\sqrt{-g}g^{0i}\right\}\partial_{0}+\frac{1}{g^{00}\sqrt{-g}}\partial_{i}\sqrt{-g}g^{ij}\partial_{j}+\frac{m^{2}-\lambda R}{g^{00}}\right)\psi=0.$ (15)
[…]
$\mathcal{D}_{\mu}\left(\sqrt{-g}g^{\mu\nu}\mathcal{D}_{\nu}\right)$ $=\mathcal{D}_{0}\left(\sqrt{-g}g^{00}\mathcal{D}_{0}\right)+\mathcal{D}_{0}\left(\sqrt{-g}g^{01}\mathcal{D}_{1}\right)+\mathcal{D}_{0}\left(\sqrt{-g}g^{02}\mathcal{D}_{2}\right)+\mathcal{D}_{0}\left(\sqrt{-g}g^{03}\mathcal{D}_{3}\right)$
$+\mathcal{D}_{1}\left(\sqrt{-g}g^{10}\mathcal{D}_{0}\right)+\mathcal{D}_{1}\left(\sqrt{-g}g^{11}\mathcal{D}_{1}\right)+\mathcal{D}_{1}\left(\sqrt{-g}g^{12}\mathcal{D}_{2}\right)+\mathcal{D}_{1}\left(\sqrt{-g}g^{13}\mathcal{D}_{3}\right)$
$+\mathcal{D}_{2}\left(\sqrt{-g}g^{20}\mathcal{D}_{0}\right)+\mathcal{D}_{2}\left(\sqrt{-g}g^{21}\mathcal{D}_{1}\right)+\mathcal{D}_{2}\left(\sqrt{-g}g^{22}\mathcal{D}_{2}\right)+\mathcal{D}_{2}\left(\sqrt{-g}g^{23}\mathcal{D}_{3}\right)$
$+\mathcal{D}_{3}\left(\sqrt{-g}g^{30}\mathcal{D}_{0}\right)+\mathcal{D}_{3}\left(\sqrt{-g}g^{31}\mathcal{D}_{1}\right)+\mathcal{D}_{3}\left(\sqrt{-g}g^{32}\mathcal{D}_{2}\right)+\mathcal{D}_{3}\left(\sqrt{-g}g^{33}\mathcal{D}_{3}\right)$ (16)
So, divided by $g^{00}$ we obtain
$\left(D_{0}^{2}+\frac{1}{g^{00}\sqrt{-g}}\left\{D_{i},\sqrt{-g}g^{0i}\right\}D_{0}+\frac{1}{g^{00}\sqrt{-g}}D_{i}\sqrt{-g}g^{ij}D_{j}+\frac{\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}}{g^{00}}+\frac{m^{2}-\frac{R}{4}}{g^{00}}\right)\psi=0.$ (17)
Which is a direct generalization of Eq. (15).
To obtain the Hamiltonian for the given generalized Klein-Gordon equation and its formulation, we proceed step by step using the Feshbach-Villars (FV) transformation and the framework provided in the question: The generalized Klein-Gordon equation in covariant form is given as:
$\left(D_{0}^{2}+\frac{1}{g^{00}\sqrt{-g}}\left\{D_{i},\sqrt{-g}g^{0i}\right\}D_{0}+\frac{1}{g^{00}\sqrt{-g}}D_{i}\sqrt{-g}g^{ij}D_{j}+\frac{\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}}{g^{00}}+\frac{m^{2}-\frac{R}{4}}{g^{00}}\right)\psi=0,$ (18)
where $D_{\mu}=\partial_{\mu}+\Gamma_{\mu}$ represents the covariant derivative.
The FV transformation is applied to reformulate the wave function $\psi$ in terms of two components $\phi$ and $\chi$:
$\psi=\phi+\chi,\quad i\left(D_{0}+\mathrm{Y}\right)\psi=m(\phi-\chi),$ (19)
where:
$\mathrm{Y}=\frac{1}{2g^{00}\sqrt{-g}}\left\{D_{i},\sqrt{-g}g^{0i}\right\},$ (20)
Using the FV transformation, the Klein-Gordon equation is rewritten as a first-order equation for $\psi$.
Substituting the definitions of $\phi$ and $\chi$, we derive a coupled system of equations:
$i\partial_{0}\psi=\mathcal{H}\psi,$ (21)
where the Hamiltonian operator $\mathcal{H}$ in terms of the components becomes:
$\mathcal{H}=\begin{pmatrix}\mathrm{Y}+\frac{m^{2}-\frac{R}{4}}{2m}+\frac{1}{4}\omega_{0}^{ab}\gamma_{ab}-iqA_{0}&-\frac{1}{2m}\left(\frac{1}{\sqrt{-g}}D_{i}\sqrt{-g}g^{ij}D_{j}+\frac{\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}}{g^{00}}\right)\\ -\frac{1}{2m}\left(\frac{1}{\sqrt{-g}}D_{i}\sqrt{-g}g^{ij}D_{j}+\frac{\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}}{g^{00}}\right)&-\mathrm{Y}-\frac{m^{2}-\frac{R}{4}}{2m}+\frac{1}{4}\omega_{0}^{ab}\gamma_{ab}-iqA_{0}\end{pmatrix}.$ (22)
For practical cases, simplifications often involve: (i) Assuming specific forms of the metric $g_{\mu\nu}$, (ii) Neglecting terms like $\frac{R}{4}$ or $\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}$ in weak field approximations, and (iii) Approximating $\mathrm{Y}$ in terms of first-order derivatives of the metric and connections. The Hamiltonian is now expressed in a block-matrix form suitable for analysis of dynamics. The FV transformation splits $\psi$ into positive and negative energy states $(\phi,\chi)$.
To rewrite the Hamiltonian in terms of Pauli matrices, we use the Feshbach-Villars formalism and express the two-component wave function $\psi=(\phi,\chi)^{T}$. The generalized Klein-Gordon Hamiltonian takes the form:
$\mathcal{H}=\sigma_{z}\left(\mathrm{Y}+\frac{m^{2}-\frac{R}{4}}{2m}+\frac{1}{4}\omega_{0}^{ab}\gamma_{ab}-iqA_{0}\right)+\sigma_{x}\mathcal{O},$ (23)
where $\sigma_{x},\sigma_{y}$, and $\sigma_{z}$ are Pauli matrices, and the operators $\mathrm{Y}$ and $\mathcal{O}$ are defined as follows: Operator $\mathrm{Y}$:
$\mathrm{Y}=\frac{1}{2g^{00}\sqrt{-g}}\left\{D_{i},\sqrt{-g}g^{0i}\right\},$ (24)
which includes contributions from the metric and the connections. Operator $\mathcal{O}$:
$\mathcal{O}=\frac{1}{2m}\left(\frac{1}{\sqrt{-g}}D_{i}\sqrt{-g}g^{ij}D_{j}+\frac{\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}}{g^{00}}\right),$ (25)
which accounts for the kinetic terms and electromagnetic interactions.
III.1 (1+2) dimension
The term $\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}$ represents the interaction between the electromagnetic field tensor $F_{\mu\nu}$ and the spinor field via the spin matrix $\sigma^{\mu\nu}$. Here’s how to evaluate it: The field strength tensor is defined as:
$F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu},$ (26)
where $A_{\mu}$ is the electromagnetic four-potential. For a given electromagnetic configuration: $F_{0i}$: Electric field components $(E_{i})$. $F_{ij}$: Magnetic field components $(B_{k})$, with $\epsilon^{ijk}B_{k}=F_{ij}$.
In $(1+2)$-dimensional spacetime, $F_{\mu\nu}$ simplifies due to fewer spatial dimensions. The spin matrix for a spinor field is defined as:
$\sigma^{\mu\nu}=\frac{i}{2}\left[\gamma^{\mu},\gamma^{\nu}\right],$ (27)
where $\gamma^{\mu}$ are the gamma matrices satisfying the Clifford algebra:
$\{\gamma^{\mu},\gamma^{\nu}\}=2g^{\mu\nu}.$ (28)
In $(1+2)$-dimensional spacetime, the gamma matrices are typically:
$\gamma^{0}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\quad\gamma^{1}=\begin{pmatrix}0&\sigma^{1}\\ \sigma^{1}&0\end{pmatrix},\quad\gamma^{2}=\begin{pmatrix}0&\sigma^{2}\\ \sigma^{2}&0\end{pmatrix},$ (29)
where $\sigma^{1}$ and $\sigma^{2}$ are the standard Pauli matrices. The commutator produces:
$\sigma^{\mu\nu}=\frac{i}{2}\begin{pmatrix}[\gamma^{\mu},\gamma^{\nu}]\end{pmatrix}.$ (30)
Using the decomposition of $F_{\mu\nu}$ into electric $(E_{i})$ and magnetic $(B_{i})$ fields:
$\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}=e\gamma^{0}\vec{\gamma}\cdot\vec{E}-e\vec{\Sigma}\cdot\vec{B},$ (31)
where: $\vec{E}$ is the electric field, $\vec{B}$ is the magnetic field and $\vec{\Sigma}$ is the spin operator.
For a purely magnetic field in $(1+2)$, with $F_{12}=B$, the spin interaction term reduces to:
$\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}=eB\sigma^{12}.$ (32)
If $\sigma^{12}$ corresponds to a specific representation of the spinor field in $(1+2)$-dimensional spacetime, this term captures the coupling of the spin to the magnetic field.
For a purely electric field in $(1+2)$, with $F_{01}=E_{x}andF_{02}=E_{y}$, the term becomes:
$\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}=e(\gamma^{0}\gamma^{1}E_{x}+\gamma^{0}\gamma^{2}E_{y}).$ (33)
III.2 in (1+3)-dimensions
To evaluate $\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}$ in $(1+3)$-dimensional spacetime, we need to compute the interaction between the electromagnetic field tensor $F_{\mu\nu}$ and the spin matrices $\sigma^{\mu\nu}.$
The electromagnetic field tensor in $(1+3)$-dimensions is:
$F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu},$ (34)
where $A_{\mu}=(A_{0},\vec{A})$ is the four-potential: Components of $F_{\mu\nu}$ are $F_{0i}=E_{i}$: Electric field components and $F_{ij}=\epsilon_{ijk}B_{k}$: Magnetic field components.
Explicitly:
$F_{\mu\nu}=\begin{pmatrix}0&-E_{x}&-E_{y}&-E_{z}\\ E_{x}&0&-B_{z}&B_{y}\\ E_{y}&B_{z}&0&-B_{x}\\ E_{z}&-B_{y}&B_{x}&0\end{pmatrix}.$ (35)
The spin matrix is:
$\sigma^{\mu\nu}=\frac{i}{2}\left[\gamma^{\mu},\gamma^{\nu}\right],$ (36)
where $\gamma^{\mu}$ are the gamma matrices satisfying the Clifford algebra:
$\{\gamma^{\mu},\gamma^{\nu}\}=2g^{\mu\nu}.$ (37)
In the Dirac representation:
$\gamma^{0}=\begin{pmatrix}I&0\\ 0&-I\end{pmatrix},\quad\gamma^{i}=\begin{pmatrix}0&\sigma^{i}\\ -\sigma^{i}&0\end{pmatrix},$ (38)
where $\sigma^{i}$ are the Pauli matrices. The commutators yield:
$\sigma^{\mu\nu}=\begin{cases}\sigma^{0i}=i\gamma^{0}\gamma^{i},&\text{time-space components,}\\ \sigma^{ij}=-\epsilon^{ijk}\begin{pmatrix}\sigma^{k}&0\\ 0&\sigma^{k}\end{pmatrix},&\text{space-space components.}\end{cases}$
Substituting $F_{\mu\nu}\text{and}\sigma^{\mu\nu}$:
$\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}=e\gamma^{0}\vec{\gamma}\cdot\vec{E}-e\vec{\Sigma}\cdot\vec{B}.$ (39)
With
$\vec{E}=(E_{x},E_{y},E_{z}),\quad\vec{\gamma}=(\gamma^{1},\gamma^{2},\gamma^{3}).$ (40)
The electric field term is:
$e\gamma^{0}\vec{\gamma}\cdot\vec{E}=e\gamma^{0}(\gamma^{1}E_{x}+\gamma^{2}E_{y}+\gamma^{3}E_{z}).$ (41)
Now, where
$\vec{B}=(B_{x},B_{y},B_{z}),\quad\vec{\Sigma}=\frac{i}{2}\epsilon_{ijk}\sigma^{ij}.$ (42)
The magnetic field term is:
$-e\vec{\Sigma}\cdot\vec{B}=-e\begin{pmatrix}\vec{\sigma}&0\\ 0&\vec{\sigma}\end{pmatrix}\cdot\vec{B}.$ (43)
Combining both terms:
$\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}=e\gamma^{0}(\gamma^{1}E_{x}+\gamma^{2}E_{y}+\gamma^{3}E_{z})-e\begin{pmatrix}\vec{\sigma}\cdot\vec{B}&0\\ 0&\vec{\sigma}\cdot\vec{B}\end{pmatrix}.$ (44)
III.2.1 Special Case: Uniform Fields
  • Uniform Magnetic Field $(\vec{E}=0)$
    $\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}=-e\vec{\Sigma}\cdot\vec{B}.$ (45)
  • Uniform Electric Field $(\vec{B}=0)$
    $\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}=e\gamma^{0}\vec{\gamma}\cdot\vec{E}.$ (46)
IV Conclusion
The Feshbach-Villars formalism for a spin-1/2 particle in curved spacetime provides a robust method for exploring relativistic quantum mechanics in the presence of gravitational and electromagnetic fields. By deriving the Hamiltonian form of the Dirac equation and applying the FV transformation, we have developed a two-component formulation that clearly separates positive and negative energy states. This formalism presents several advantages:
  • It allows a more intuitive interpretation of the wavefunction components with respect to particle and antiparticle contributions.
  • The resulting Hamiltonian, expressed in terms of Pauli matrices, facilitates the analysis of spin-dependent effects in curved spacetime.
  • The approach naturally incorporates both gravitational and electromagnetic interactions, making it suitable for exploring various physical scenarios.
Our analysis in both (1+2) and (1+3) dimensions showcases the versatility of the formalism and highlights the specific forms of spin-field interactions across different spacetime configurations. The explicit treatment of the term involving the electromagnetic field tensor and spin matrices provides a clear understanding of how spin couples to external fields in curved spacetime. This work lays the groundwork for future studies on relativistic quantum effects in strong gravitational fields, the behavior of spin in curved spacetime, and potential applications in areas such as early universe cosmology and the physics of compact objects. Further research could investigate the implications of this formalism for quantum field theory in curved spacetime and explore possible experimental tests of the predicted effects.
?refname?
  • 1 (2025) H. Feshbach and F. M. H. Villars, Rev. Modern Phys. 30, 24 (1958).
  • 2 (2025) B. A. Robson and D. S. Staudte, J. Phys. A : Math. Gen 29, 157 (1996).
  • 3 (2025) D. S. Staudte, J. Phys. A 29, 169 (1996).
  • 4 (2025) M. Merad, L. Chetouani, and A. Bounames, Phys. Lett. A 267, 225 (2000).
  • 5 (2025) A. Bounames and L. Chetouani, Phys. Lett. A 279, 139 (2001).
  • 6 (2025) S. Haouat and L. Chetouani, Eur. Phys. J. C 41, 297 (2005).
  • 7 (2025) B Guettou1 and L Chetouani, Phys. Scr. 74, 12–18 (2006).
  • 8 (2025) N. Brown, Z. Papp, and R. M. Woodhouse, Few-Body Systems 57, 103 (2015).
  • 9 (2025) B. Motamedi, T. Shannon, and Z. Papp, Few-Body Systems 60, 65 (2019).
  • 10 (2025) D. Wingard, A. Garcia Vallejo and Z. Papp, Few-Body Syst 65, 30 (2024).
  • 11 (2025) A. J. Silenko, Phys. Rev. A 77, 012116 (2008).
  • 12 (2025) A. J. Silenko, Phys. Rev. D 88, 045004 (2013).
  • 13 (2025) A. Bouzenada, A. Boumali, E.O. Silva, Ann. Phys. 458, 169479 (2023)
  • 14 (2025) A. Bouzenada, A. Boumali, R. Vitória, F. Ahmed, M. Al-Raeei, Nucl. Phys. B 994, 116288 (2023)
  • 15 (2025) A. Bouzenada, A. Boumali, F. Ahmed, Nucl. Phys. B 1007, 116682 (2024)
  • 16 (2025) A.Bouzenada and A.Boumali. Ann. Physics 27, 169302 (2023).
  • 17 (2025) E. Schrodinger, Sitz. Preuss. Akad. Wiss. Berlin 105 (1932).
Appendix A Proof of Eq. (13)
The Dirac equation in curved spacetime is:
$\left[i\gamma^{\mu}\mathcal{D}_{\mu}-m\right]\Psi=0.$ (47)
Squaring the operator using:
$\left[i\gamma^{\mu}\mathcal{D}_{\mu}-m\right]\left[i\gamma^{\mu}\mathcal{D}_{\mu}+m\right],$ (48)
we get:
$\left[i\gamma^{\mu}\mathcal{D}_{\mu}-m\right]\left[i\gamma^{\mu}\mathcal{D}_{\mu}+m\right]=\left(i\gamma^{\mu}\mathcal{D}_{\mu}\right)^{2}-m^{2}.$ (49)
Expand $(i\gamma^{\mu}\mathcal{D}_{\mu})^{2}$:
$\left(i\gamma^{\mu}\mathcal{D}_{\mu}\right)^{2}=-\gamma^{\mu}\gamma^{\nu}\mathcal{D}_{\mu}\mathcal{D}_{\nu}.$ (50)
The gamma matrices satisfy the Clifford algebra:
$\gamma^{\mu}\gamma^{\nu}=g^{\mu\nu}+\sigma^{\mu\nu},\quad\text{where }\sigma^{\mu\nu}=\frac{1}{2}\left[\gamma^{\mu},\gamma^{\nu}\right].$ (51)
Substitute:
$-\gamma^{\mu}\gamma^{\nu}\mathcal{D}_{\mu}\mathcal{D}_{\nu}=-g^{\mu\nu}\mathcal{D}_{\mu}\mathcal{D}_{\nu}-\sigma^{\mu\nu}\mathcal{D}_{\mu}\mathcal{D}_{\nu}.$ (52)
The covariant derivative $\mathcal{D}_{\mu}=\partial_{\mu}+\Gamma_{\mu}+\text{spin connection}$ satisfies:
$[\mathcal{D}_{\mu},\mathcal{D}_{\nu}]=\frac{1}{4}R_{\mu\nu\lambda\sigma}\gamma^{\lambda\sigma}-iqF_{\mu\nu}.$ (53)
From this:
$g^{\mu\nu}\mathcal{D}_{\mu}\mathcal{D}_{\nu}=\frac{1}{\sqrt{-g}}\mathcal{D}_{\mu}\left(\sqrt{-g}g^{\mu\nu}\mathcal{D}_{\nu}\right)-\frac{1}{4}R,$ (54)
where $R$ is the Ricci scalar.
For the spin terms:
$\sigma^{\mu\nu}F_{\mu\nu}=\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}.$ (55)
Substituting back:
$\left(i\gamma^{\mu}\mathcal{D}_{\mu}\right)^{2}=-\left(\frac{1}{\sqrt{-g}}\mathcal{D}_{\mu}\left(\sqrt{-g}g^{\mu\nu}\mathcal{D}_{\nu}\right)-\frac{1}{4}R+\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}\right).$ (56)
Adding the mass term $-m^{2}$:
$\left[i\gamma^{\mu}\mathcal{D}_{\mu}-m\right]\left[i\gamma^{\mu}\mathcal{D}_{\mu}+m\right]\Psi=\left(-\frac{1}{\sqrt{-g}}\mathcal{D}_{\mu}\left(\sqrt{-g}g^{\mu\nu}\mathcal{D}_{\nu}\right)+\frac{1}{4}R-\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}+m^{2}\right)\Psi=0.$ (57)
The Klein-Gordon equation derived from the squared Dirac operator is:
$\left(\frac{1}{\sqrt{-g}}\mathcal{D}_{\mu}\left(\sqrt{-g}g^{\mu\nu}\mathcal{D}_{\nu}\right)-\frac{1}{4}R+\frac{ie}{2}F_{\mu\nu}\sigma^{\mu\nu}+m^{2}\right)\Psi=0.$ (58)
Here, the derivation is exact under the assumptions of: Minimal coupling to gravity and electromagnetism, Metric compatibility $(\nabla_{\mu}g_{\alpha\beta}=0)$ and Torsion-free spacetime. It is important to mention that (58) was originally derived by Schrodinger (17, 2025).